A question dealing with solving functions says, Find the indicated values, where g(t)=t^2-t and f(x)=1+x. They give you this formula, f(2g(1)).
Answer Choices are: 1, 3, 30, -1

Answer by rothauserc(4718)

g(1) = 1^2 - 1 = 0
:
2g(1) = 2 * 0 = 0
:
f(0) = 1 + 0 = 1
:
therefore,
:
************
f(2g(1)) = 1
************
: